1. In analysis of variance, MS between-groups provides a measure of

  the total variability for the set of N scores, standard deviation, the overall mean for the set of N scores, differences among treatment means

2. An independent-measures research study compares three treatment conditions using a sample of n = 5 in each treatment. For this study, the three totals are, T1 = 5, T2 = 10, and T3 = 15. What value would be obtained for the grand mean, Mgrand?

2,10,15,5
Answer #1

(a) In Anova the MS between groups provides a measure of variance in difference among treatment means.

(2) Grand Mean = (5 + 10 + 15) / 15 = 30 / 15 = 2